All members have constant El B.M. to be evaluated in terms of w or PL, H solve by consistent deformation 8t 7 m 9m

All members have constant El B.M. to be evaluated in terms of w or PL, H solve by consistent deformation 6m 8m
Answer #1

Given atim am ym ym DTTI ATU حي 1.8m 1.8m of the Given structure The static indeterminacy No of unknown Reaction - 3 is 6-3 -Consider Section at a between distance from 2 tim a -X x varies from 0 to 7 m. 82=1 unit 33=2 unit = 1 unit here, My = 0 KN-m3. Here, Ma? 2 3.24+x2 1 + 2x1.8 2 + 2x x3 2 ma 2 m2 - 7 tom mz 1 t-m. distance at a section a Now Consider section from Bi nM3 Ma m2 mi - Member Origin Length 1 D o-7 -7 1 0-9 3.24+x² x С CB -(7x) 1 07 91 9 BA 12 Now, the flexiblity matrix F da. mi.F23 - For x.da t a da GG EI 7 9x a3 L - 1035 EI LEI 7 F22 x2 da t 49 da + (7x)? da. СТ СТ СТ 7 + 49847 + Í (70 BEI 3 - 669.66103.5 810 [f] = 504 669-66 12 -504 EI 23 103.5 -112 Now, the matrix due to actual loading Say COOL] DoLi - Ma. Mi. da. I 1 +[Dec] = 1 ET 6874-47 3889.62 839.16 at the There no deflection of Rotation Joint Di say cos Dis place Joint matrix [DS] = O [3 1035 {F} = 810 soy -504 669-66 -112 -112 EL 103.5 23 2858 18 - {ady f] - -12861.81 38556 28840 86 7917.75 38556 EI) 12861.8- 8.968t (1) O = 1.567t (e) 3.73 ton (5) Og - Net vertical force + Vp 2x 12.6 % 25.2- 8.968 V = 16.232 t (1) . Net horizontalTo solve using consistent deformation method, at first the static indeterminacy is determined. After that the flexiblity matrix and deflection matrix is developed. Hence, the compatiblity equation is used to get the unknowns.

Note:- If in any kind of doubt, please mention in comment.

Thanks!

Feedback appreciated!